# Boson-fermion pairing and condensation in two-dimensional Bose-Fermi mixtures

### v2

@@ 4.5 Tan’s contact parameter and composite-fermion density
The presence of $n_{\mathrm{CF}}$ in Eq. (59) suggests comparing $n_{\mathrm{CF}}$ defined by Eq. (33) to the expression ${C_{\mathrm{BF}}}/(8\pi m_{r}{\varepsilon_{0}}{n_{\mathrm{F}}})$, in order to see when deviations from the asymptotic expression occur. This is done in Fig. 15(b) for several values of the concentration $x$ (full lines and dotted lines, respectively). One sees that deviations are most significant at small concentrations, but for $g\gtrsim 2.5$ the contact essentially coincides with its asymptotic expression for all cases. We notice furthermore that for $g\gtrsim 1$, the quantity $n_{\mathrm{CF}}$ essentially coincides with the number of bosons out of the condensate, $n_{\mathrm{B}}-n_{0}$, (represented by the dashed lines in Fig. 15(b)). This justifies the assumption ${n_{\mathrm{CF}}}\simeq{n_{\mathrm{B}}}-n_{0}$ which we made in Sec. 3.4 when discussing the strong-coupling limit of our theory.
@@ 4.6 Identification of different physical regions in the coupling-concentration plane
Based on the results obtained in Secs. 3.23.44.3, and 4.4, the coupling ($g=-\ln(k_{\rm F}a_{\rm BF})$) vs. boson concentration ($x=n_{\rm B}/n_{\rm F}$) plane of a BF mixture can be divided in different physical region.
@@
Specifically, in Sec. 3.2 and  3.4 two different regimes were identified when varying the bosonic concentration for a fixed strong BF attraction. Analogously, two different regimes were identified in Sec. 4.3 when varying the coupling strength, while the analysis of the condensate fraction of Sec. 4.4 added further insight into the evolution with the coupling strength.
@@ 5 Conclusions and outlook
Similarly to the 3D case, we uncovered a nearly universal behavior for the condensate fraction and bosonic momentum distribution with respect to the boson concentration. Contrary to the 3D case, however, we have found that the universal condensate fraction does not match the quasiparticle residue $Z$ of the Fermi polaron. It would be interesting to understand whether the equivalence between these two quantities observed in 3D has an intrinsic explanation or is simply an accidental degeneracy that is removed in 2D.
A further difference with respect to the 3D case is the hybrid nature of the composite fermions that form for sufficiently strong BF attraction. Rather unusual features are found within the momentum distribution function, except for the case $x=1$ in which the mixture “fermionizes” completely in a gas of non-interacting point-like fermionic molecules, similarly to what is obtained in 3D for all $x\leq 1$. The origin of this difference for $x<1$ is in the residual condensate density $n_{0}$ which, albeit small, proves sufficient to generate a significant hybridization through the hybridization energy-scale $\Delta_{0}\propto(n_{0}{\varepsilon_{0}})^{1/2}$.
A discussion on possible refinements of the present approach is finally in order. The absence of a quantum phase transition for the condensate fraction found in the present work recalls the absence of the polaron-to-molecule transition that is found in the single impurity limit of a Fermi polaron when the polaron is described by a non-self-consistent $T$-matrix self-energy [97]. In this limit, it has been pointed out that a correct description of the atom-dimer interaction is crucial in bringing about this transition [97, 98].
@@
Insight from the polaron limit may suggest that the absence of a critical coupling strength in 2D is simply an artifact of the approximation for the self-energy here adopted. We do not exclude such a possibility, in particular at small values of $x$ which are closer to the single impurity limit. We however stress that, in experiments, the exponential vanishing of the condensate at large coupling strength, which we have found in the present work, would hardly be distinguishable from an exact vanishing of the condensate (leaving aside the fact that at finite temperature, as experimentally relevant, one should consider a quasi-condensate in 2D, introducing a further source of smearing of the transition).
@@
In addition, we point out that, since the polaron-to-molecule quantum phase transition is proven to occur specifically in the impurity limit, one cannot exclude that in the high concentration regime, where the polaron picture breaks down, this transition could be suppressed or significantly modified by quantum fluctuations. It is known that in 3D, with a finite density of impurities, thermal and quantum fluctuations smoothen the discontinuity arising in the single impurity limit thus shifting the critical coupling to higher values [127, 33]. An analogous and stronger role may be played by quantum fluctuations in 2D. To settle this point, a detailed calculation of the three-body diagrammatic contributions for finite densities beyond the polaronic regime should be addressed in future work.
It is also worth discussing the relevance of the present investigation to BF mixtures realized with transition metal dichalcogenides. The first issue is whether the effective interaction between excitons and excess charge carriers (the bosons and fermions of the mixture, respectively) can be modeled by a contact potential of the type utilized in the present work. In TMD systems, excitons and charge carriers can bind into a molecular trion state, which controls the effective BF interaction. The binding energy of the trion is typically at least one order of magnitude smaller than the exciton binding energy [128], and of the order of the Fermi energy of the excess charge carrier. Under these conditions, it has been shown that the effective BF interaction can indeed be reasonably modeled by a contact interaction [129]. A difference with respect to the model discussed in the present paper may instead be represented by the presence of Coulomb interactions between fermions, in particular intra-species interactions between excess charge carriers. It can, however, be argued that [130], provided the doping is not too low [90, 56] (in order to be away from the regime of Wigner crystallization and reduced screening), their effect will be mostly to renormalize the single-particle dispersions of excess charges, an effect that can be safely described by the parameters of Fermi liquid theory (notably, by the effective mass of the charges). Concerning, finally, exciton-exciton interaction, it is expected to be small due to the tightly-bound nature of the excitons in TMDs that makes their dipole moment small, with a sign of the interaction that depends on the orientation of the dipoles. We have already mentioned that the mechanical stability of BF mixtures requires a repulsive interaction. In this respect, bilayer TMDs with interlayer excitons would be preferable, since in this case the dipole moments will be essentially parallel to each other (we notice that the distance between layers[42] is comparable with the exciton radius [90]), thus ensuring the interaction to be repulsive. It is however not clear if such a repulsion would be sufficient to guarantee stability.
@@
All numerical data necessary to reproduce Figs. 3-15 are available online [131].
@@ Acknowledgements
L.P.  and P.P.  acknowledge financial support from the Italian Ministry of University and Research (MUR) under project PRIN2022, Contract No. 2022523NA7. P.P. also acknowledges financial support from MUR project PE0000023-NQSTI (Italy) and from European Union - NextGenerationEU through the Italian Ministry of University and Research (MUR) under PNRR - M4C2 - I1.4 Contract No. CN00000013.
@@ Appendix A Particle-particle ladder $\Gamma$(${\mathbf{P}}$, $\Omega$)
$\Gamma\left({\mathbf{P}},\Omega\right)=\frac{2\pi/m_{r}}{\ln\left(\varepsilon_{0}/\mathcal{A}\right)+i\pi\textrm{sgn}\left(\textrm{Im}\left[\mathcal{A}\right]\right)}.$ (A.9)
@@ D.2 Fermion momentum distribution for small hybridization energy $\Delta_{0}$
$\frac{\Delta_{\mathrm{CF}}^{2}}{i\omega+\xi^{\mathrm{B}}_{{\mathbf{k}}}+\mu_{\mathrm{CF}}}\approx-\frac{\Delta_{\mathrm{CF}}^{2}}{\mu_{\mathrm{B}}}={2\pi n_{\mathrm{CF}}\over m_{r}}\equiv\Sigma_{\mathrm{F}}^{0}$ (D.10)

### v3

@@ 4.5 Tan’s contact parameter and composite-fermion density
The presence of $n_{\mathrm{CF}}$ in Eq. (63) suggests comparing $n_{\mathrm{CF}}$ defined by Eq. (33) to the expression ${C_{\mathrm{BF}}}/(8\pi m_{r}{\varepsilon_{0}}{n_{\mathrm{F}}})$, in order to see when deviations from the asymptotic expression occur. This is done in Fig. 17(b) for several values of the concentration $x$ (full lines and dotted lines, respectively). The deviations are most significant at small concentrations, but for $g\gtrsim 2.5$ the contact essentially coincides with its asymptotic expression for all cases. We also notice that for $g\gtrsim 1$, the quantity $n_{\mathrm{CF}}$ essentially coincides with the number of bosons out of the condensate, $n_{\mathrm{B}}-n_{0}$, (represented by the dashed lines in Fig. 17(b)). This justifies the assumption ${n_{\mathrm{CF}}}\simeq{n_{\mathrm{B}}}-n_{0}$ which we made in Sec. 3.4 when discussing the strong-coupling limit of our theory.
@@ 4.6 Identification of different physical regions in the coupling-concentration plane
Based on the results obtained in Secs. 3.23.44.3, and 4.4, the coupling ($g=-\ln(k_{\rm F}a_{\rm BF})$) vs. boson concentration ($x=n_{\rm B}/n_{\rm F}$) plane of a BF mixture can be divided into different physical regions.
@@
Specifically, in Sec. 3.2 and  3.4 two different regimes were identified when varying the bosonic concentration for a fixed strong BF attraction. Similarly, two different regimes were identified in Sec. 4.3 when varying the coupling strength, while the analysis of the condensate fraction of Sec. 4.4 added further insight into the evolution with the coupling strength.
@@ 5 Conclusions and outlook
Similarly to the 3D case, we uncovered a nearly universal behavior for the condensate fraction and bosonic momentum distribution with respect to the boson concentration. However, in contrast to the 3D case, we have found that the universal condensate fraction does not match the quasiparticle residue $Z$ of the Fermi polaron. It would be interesting to understand whether the equivalence between these two quantities observed in 3D has an intrinsic explanation or is simply an accidental degeneracy that is removed in 2D.
A further difference with respect to the 3D case is the hybrid nature of the composite fermions that form for sufficiently strong BF attraction. Rather unusual features are found within the momentum distribution function, except for the case $x=1$ in which the mixture “fermionizes” completely in a gas of non-interacting point-like fermionic molecules, similarly to what is obtained in 3D for all $x\leq 1$. The origin of this difference for $x<1$ is in the residual condensate density $n_{0}$ that, although small, proves sufficient to generate a significant hybridization through the hybridization energy-scale $\Delta_{0}\propto(n_{0}{\varepsilon_{0}})^{1/2}$.
Finally, a discussion on possible refinements of the present approach is in order. The absence of a quantum phase transition for the condensate fraction found in the present work recalls the absence of the polaron-to-molecule transition that is found in the single impurity limit of a Fermi polaron when the polaron is described by a non-self-consistent $T$-matrix self-energy [97]. In this limit, it has been pointed out that a correct description of the atom-dimer interaction is crucial in bringing about this transition [97, 98].
@@
Insight from the polaron limit may suggest that the absence of a critical coupling strength in 2D is simply an artifact of the approximation for the self-energy here adopted. We do not exclude such a possibility, in particular at small values of $x$ which are closer to the single impurity limit. We, however, stress that, in experiments, the exponential vanishing of the condensate at large coupling strength, which we have found in the present work, would hardly be distinguishable from an exact vanishing of the condensate (leaving aside the fact that at finite temperature, as experimentally relevant, one should consider a quasi-condensate in 2D, introducing a further source of smearing of the transition).
@@
In addition, we point out that, since the polaron-to-molecule quantum phase transition is proven to occur specifically in the impurity limit, one cannot exclude that in the high-concentration regime, where the polaron picture breaks down, this transition could be suppressed or significantly modified by quantum fluctuations. It is known that in 3D, with a finite density of impurities, thermal and quantum fluctuations smoothen the discontinuity arising in the single impurity limit thus shifting the critical coupling to higher values [131, 33]. An analogous and stronger role may be played by quantum fluctuations in 2D. To settle this point, a detailed calculation of the three-body diagrammatic contributions for finite densities beyond the polaronic regime should be addressed in future work.
It is also worth discussing the relevance of the present investigation to BF mixtures realized with transition metal dichalcogenides. The first issue is whether the effective interaction between excitons and excess charge carriers (the bosons and fermions of the mixture, respectively) can be modeled by a contact potential of the type utilized in the present work. In TMD systems, excitons and charge carriers can bind into a molecular trion state, which controls the effective BF interaction. The binding energy of the trion is typically at least one order of magnitude smaller than the exciton binding energy [132], and of the order of the Fermi energy of the excess charge carrier. Under these conditions, it has been shown that the effective BF interaction can indeed be reasonably modeled by a contact interaction [133]. A difference with respect to the model discussed in the present paper may instead be represented by the presence of Coulomb interactions between fermions, in particular intra-species interactions between excess charge carriers. However, it can be argued that [134], provided the doping is not too low [90, 56] (in order to be away from the regime of Wigner crystallization and reduced screening), their effect will mostly be to renormalize the single-particle dispersions of excess charges, an effect that can be safely described by the parameters of Fermi liquid theory (notably, by the effective mass of the charges). Concerning, finally, exciton-exciton interaction, it is expected to be small because of the tightly-bound nature of the excitons in TMDs that makes their dipole moment small, with a sign of the interaction that depends on the orientation of the dipoles. We have already mentioned that the mechanical stability of BF mixtures requires a repulsive interaction. In this respect, bilayer TMDs with interlayer excitons would be preferable, since in this case the dipole moments will be essentially parallel to each other (we notice that the distance between layers[42] is comparable with the exciton radius [90]), thus ensuring that the interaction isrepulsive. However, it is not clear if such a repulsion would be sufficient to guarantee stability.
@@
All the numerical data necessary to reproduce Figs. 3-17 are available online [135].
@@ Acknowledgements
L.P.  and P.P.  acknowledge financial support from the Italian Ministry of University and Research (MUR) under project PRIN2022, Contract No. 2022523NA7. P.P. also acknowledges financial support from MUR project PE0000023-NQSTI (Italy) financed by the European Union – Next Generation EU and from European Union - NextGeneration EU through MUR under PNRR - M4C2 - I1.4 Contract No. CN00000013.
@@ Appendix A Particle-particle ladder $\Gamma$(${\mathbf{P}}$, $\Omega$)
$\Gamma\left({\mathbf{P}},\Omega\right)=\frac{2\pi/m_{r}}{\ln\left(\varepsilon_{0}/\mathcal{A}\right)+i\pi\textrm{sgn}(\Omega)}.$ (A.9)
@@ D.2 Fermion momentum distribution for small hybridization energy $\Delta_{0}$
$\frac{\Delta_{\mathrm{CF}}^{2}}{i\omega+\xi^{\mathrm{B}}_{{\mathbf{k}}}}\approx-\frac{\Delta_{\mathrm{CF}}^{2}}{\mu_{\mathrm{B}}}={2\pi n_{\mathrm{CF}}\over m_{r}}\equiv\Sigma_{\mathrm{F}}^{0}$ (D.10)